# capbtabboxtable[][.45] (eccv) Package eccv Warning: Package ‘hyperref’ is loaded with option ‘pagebackref’, which is *not* recommended for camera-ready version

### v1

@@ Abstract
Our objective is to discover and localize monotonic temporal changes in a sequence of images. To achieve this, we exploit a simple proxy task of ordering a shuffled image sequence, with ‘time’ serving as a supervisory signal since only changes that are monotonic with time can give rise to the correct ordering. We also introduce a flexible transformer-based model for general-purpose ordering of image sequences of arbitrary length with built-in attribution maps. After training, the model successfully discovers and localizes monotonic changes while ignoring cyclic and stochastic ones. We demonstrate applications of the model in multiple video settings covering different scene and object types, discovering both object-level and environmental changes in unseen sequences. We also demonstrate that the attention-based attribution maps function as effective prompts for segmenting the changing regions, and that the learned representations can be used for downstream applications. Finally, we show that the model achieves the state of the art on standard benchmarks for ordering a set of images.
@@
Figure 1: Localizing monotonic temporal changes. Top: satellite images (ordered left to right) taken months apart, containing several changes – some are correlated monotonically with time (such as urbanization), while others are seasonal/cyclic (such as water level). Bottom: Our model’s attribution map prediction on the sequence is able to localize the regions with monotonic time-correlated changes (in green), while being invariant to the seasonal and sporadic changes. The model is trained with no manual supervision, generalises to unseen sequences (as here), and the attribution map can also be used as a prompt to obtain segmentation.
@@ 1 Introduction
To achieve this, we propose a self-supervised proxy task, with time serving as a supervisory signal: the task is simply to order a shuffled sequence of video frames. The insight is that if a model can order the video frames, it would have learned to identify relevant (monotonic temporal-correlated) cues while disregarding other changes. The trained model can then be employed for video analysis applications where the goal is to identifying changes over time, such as developments or deforestation in satellite imagery (whilst ignoring seasonal variations) (see Figure 1) or aging signs in medical images. It can also be employed for detecting and tracking monotonic object motion over time, such as shadows caused by the movement of the sun or animals moving smoothly across the scene.
@@
In order to harness the ordering proxy task, we introduce a transformer-based model that is able (i) to perform ordering on natural images sequences, and, more importantly, (ii) to provide attribution by localizing the evidence that gives rise to its prediction. Specifically, we use a DETR-like transformer encoder-decoder architecture where the queries in the decoder are cast as an ordering index. The architecture is designed to allow a attribution map to be obtained directly as part of the proxy training. Once the model has been trained for a particular setting, such as change detection in satellite image sequences, then it can generalize to unseen videos in the same setting, requiring only a forward pass to predict the localization of the monotonic temporal changes from the attribution map.
To summarize, in this paper, we make the following four contributions: (i) we introduce a new task of discovering and localizing monotonic temporal changes in image sequences, and use temporal ordering as a self-supervised proxy loss for training. (ii) we introduce a flexible transformer-based model that can be used for ordering images of different sequence lengths, and also for localizing evidence. (iii) once trained on a setting (such as satellite images), the model is able to discover the changes correlating monotonically with time in unseen image sequences in the same setting, and we demonstrate several situations where this can be applied. Finally, (iv) we demonstrate that the trained model is able to order novel image sequences surpassing the performance of previous approaches for ordering on standard benchmarks.
@@ 3.3 Training and inference
Temporal loss. Given the ground-truth order $y\in\mathbb{Z}^{Q}:y_{i}\in\{0,1,...,F-1\}$, the model can be trained via binary cross-entropy loss. Specifically, we convert the ground-truth order into a matrix where each column is a one-hot vector indicating its position, e.g. $(1,0,2)$ as $\big{(}\begin{smallmatrix}0&1&0\\ 1&0&0\\ 0&0&1\end{smallmatrix}\big{)}$. With some notation abuse, we still call this matrix $y\in\mathbb{Z}^{F\times Q}$. The forward loss is then simply the elementwise binary cross-entropy between the two matrices: $\mathcal{L}_{f}(y,\hat{y})=\frac{1}{FQ}\sum_{i\in F}\sum_{j\in Q}\text{cross-entropy}(\hat{y}_{ij},y_{ij})$.
In practice, we find that allowing reversibility in the loss aids with training, as many changes are reversible in nature without prior knowledge of the arrow of time [64] (the sequence could equally be ordered from first to last, or last to first). To allow this, we calculate the loss as the minimum of the loss for both forward and backward sequences, i.e. $\mathcal{L}_{r}=\min(\mathcal{L}_{f}(y,\hat{y}),\mathcal{L}_{f}(\text{reverse}(y),\hat{y}))$. The loss is minimized when the sequence is in the correct order, regardless of the direction.
@@ 3.4 Discussion
From localization to segmentation. While the attribution map is useful, some applications may benefit from going beyond just localization. Here, we propose two solutions. We can directly obtain segmentation at patch-level granularity by thresholding the attribution map, together with minimal post-processing namely averaging across frames and removing small contours. We note that this method is crude as it does not distinguish pixels within a patch. Alternatively, we can use the highest activation points (i.e. centre pixels of patches) as prompt for the Segment Anything model (SAM) [33], to obtain pixel-level segmentation masks. We use the pretrained SAM model without fine-tuning.
@@ 4.1 Video datasets
To leverage the inherent temporal information in videos, specifically, we sample a sequence from a video, shuffle the frames, then train the model with the original ordering as groundtruth. This approach enables the attribution map to identify temporally coherent changes that contribute to the ordering while disregarding other changes. Our study explores various sequence types across multiple domains, each with distinct cues of interest, which we summarize in Table 1 and illustrate with sample sequences in Figure 3.
@@
Timelapse scenes. We have gathered an outdoor dataset of timelapse videos from various online sources. The dataset comprises 180 static videos, and our aim is to investigate the cues that the model can extract from the scene to learn its order, as there are no specific absolute cues present.
@@ 5 Results
We present ordering results in Section 5.1. Then, as our work cuts across several domains, we consider several avenues for comparison. In Section 5.2, we compare our method with existing change detection methods, and show that discovering monotonic temporal changes is unattainable by previous methods. In Section 5.3, we compare with other self-supervised proxy tasks and show that (i) our method works better than previous works in change localization, and (ii) our method learns better representations that can be finetuned to downstream tasks. Lastly, we show in Section 5.4 that our model can serve as a general-purpose ordering method that outperforms existing methods in image ordering benchmarks.
@@ 5.1 Results on ordering video frames or image sequence
Video ordering results. The results for ordering as well as the main signals that the model can pick up, are shown by EM and EW in Table 1, with qualitative results in Figure 5. The model is able to successfully order sequences across different datasets, particularly in cases where changes are significant. It is expected that the scores are not perfect, as sampled data from videos is not guaranteed to contain ordering cues thus rendering the sequence simply unorderable, as illustrated in the Supplementary Material.
@@
Failure cases. A limitation of our model is that we do not force a one-to-one matching between queries and images, and this may result in some images being claimed by multiple different queries or by none at all, as seen in Figure 6(b). This problem can easily be resolved by allowing each image to be predicted once. However, not being able to order also provides valuable information as not all real sequences can be ordered – for example, sequences where everything is static for a period, or very stochastic. Therefore, we treat invalid orderings as a means to provide information on whether particular frames can be ordered or not.
@@ 5.2 Comparison with change detection methods
Ours 83.3 37.9 / 45.1 (SAM) 34.2 75.0
Table 2: Localization and segmentation results via the pointing game accuracy for localization, and mIoU for segmentation. The methods for segmentation (patch and SAM) are described in Section 3.4. Methods labelled as “--” are only trained on satellite images, hence do not generalize to other domains.
We compare against three other previous approaches for their ability to detect monotonic temporal changes on the Monotonic MUDS dataset, namely, a supervised Siamese Networks [26] trained on MUDS dataset for the task of urban development tracking, by highlighting the differences between buildings; and the Change Vector Analysis [42], which is a baseline for the task of change detection, highlighting all changes in an image pair without knowledge their nature. Deep Change Vector Analysis [54] is a learned version of the above CVA that has been trained specifically on satellite images.
The performance comparison is given in Table 2, and illustrated in Figure 7. As can be seen, the urban development tracking method [26] under-segments changes that are correlated with urbanisation, as it is only trained to look at buildings. The change detection methods of [42, 54] over-segments changes that are non-monotonic as it has no concept of time. Both prior methods have shortcomings in detecting urban development: the former method (and even the ground-truth provided with the original dataset) misses changes like road building and deforestation, and the latter includes many erroneous regions such as the seasonal changes in vegetation and water level. Our model is able to highlight correctly the monotonic changes while being invariant to other changes. We further note that our model discovers such changes without any prior information on what to look for. We also evaluate on two other datasets: RDS and MoCA, where we have ground-truth for the moving objects in the video; and show that we obtain favourable results.
@@
Ours 83.3 37.9 / 45.1 (SAM) 30.2
Table 3: Comparison with other self-supervised methods (left) on localization and segmentation on Monotonic MUDS. (right) on fine-tuning performance on building change detection on MUDS.
@@ 5.3 Comparison on self-supervised proxy tasks
Here, we compare to other self-supervised methods based on time. Please refer to the Supplementary Material for a discussion on the subtle differences between the proxy tasks. We include results for training baselines from scratch on MUDS, and testing on Mono-MUDS in localizing and segmenting monotonic changes. The results are shown in Table 3, where we conduct three sets of experiments, as detailed below.
@@
Third, we investigate representation learning for a target task that requires both time and localization: change detection of buildings on satellite images (like [26]). For each method, we pre-train the encoder on MUDS, freeze it, then train a lightweight head (3×3$\times$3 × deconvolutions) on top of the same dataset, and test on unseen sequences. To keep this fair, we keep the number of parameters roughly the same across methods. The results (Table 3, right column) show that our method learns good representations as compared to previous self-supervised methods in localization tasks.
@@ 5.4 Comparison with image ordering methods
We compare to two previous methods on image ordering benchmarks where the task is to arrange the images in increasing order.
Differentiable sorting networks such as DiffSort [51] and its successor DSortv2 [52] employ a parameter-free sorting network to rank scalars. We compare with these sorting models, and show that our model is capable as a general ordering network with added functionality.
Pointer Networks  [62] ranks features by using a recurrent encoder-decoder network with attention. We note that Ptr-Net is not initially designed for such a task, but for arranging a set of coordinates. We simply extend pointer networks by adding an image encoder and task the model to rank the image features from small to large. We then jointly train this encoder and the pointer network. For fair comparison, we use the same transformer encoder as our model, and use the pointer network as the decoder with similar size to our transformer decoder.

### v3

@@ Abstract
Our objective is to discover and localize monotonic temporal changes in a sequence of images. To achieve this, we exploit a simple proxy task of ordering a shuffled image sequence, with ‘time’ serving as a supervisory signal, since only changes that are monotonic with time can give rise to the correct ordering. We also introduce a transformer-based model for ordering of image sequences of arbitrary length with built-in attribution maps. After training, the model successfully discovers and localizes monotonic changes while ignoring cyclic and stochastic ones. We demonstrate applications of the model in multiple domains covering different scene and object types, discovering both object-level and environmental changes in unseen sequences. We also demonstrate that the attention-based attribution maps function as effective prompts for segmenting the changing regions, and that the learned representations can be used for downstream applications. Finally, we show that the model achieves the state-of-the-art on standard benchmarks for image ordering.
@@
Figure 1: Localizing monotonic temporal changes. Top: satellite images (ordered left to right) taken months apart, containing several changes – some are monotonic (e.g. urbanization), while others are seasonal/cyclic (e.g. water level). Bottom: Our model’s attribution map prediction on the sequence is able to localize the regions with monotonic temporal changes (in green), while being invariant to the seasonal and sporadic changes. The model is trained with no manual supervision, generalises to unseen sequences (as here), and the attribution map can also be used as a prompt to obtain segmentation.
@@ 1 Introduction
To achieve this, we propose a self-supervised proxy task, with time serving as a supervisory signal: the task is simply to order a shuffled sequence of video frames. The insight is that frames are only orderable if changes are monotonic, therefore if a model can order the video frames, it would have learned to identify relevant (monotonic temporal-correlated) cues while disregarding other changes. The trained model can then be employed for video analysis applications where the goal is to identify changes over time, such as developments or deforestation in satellite imagery (whilst ignoring seasonal variations) (see Figure 1) or aging signs in medical images. It can also be employed for detecting and tracking monotonic object motion over time, such as shadows caused by the movement of the sun or animals moving smoothly across the scene.
@@
In order to harness the ordering proxy task, we introduce a transformer-based model that is able (i) to perform ordering on natural images sequences, and, more importantly, (ii) to provide attribution by localizing the evidence that gives rise to its prediction. Specifically, we use a DETR-like transformer encoder-decoder architecture where the queries in the decoder are cast as an ordering index. The architecture is designed to allow an attribution map to be obtained directly as part of the proxy training. Once the model has been trained for a particular setting, such as change detection in satellite image sequences, then it can generalize to unseen videos in the same setting, requiring only a forward pass to predict the localization of the monotonic temporal changes from the attribution map.
To summarize, in this paper, we make the following four contributions: (i) we introduce a new task of discovering and localizing monotonic temporal changes in image sequences, and establish the link that temporal ordering can be used as a self-supervised proxy task for training; (ii) we introduce a flexible transformer-based architecture for ordering that can also automatically localize monotonic changes; (iii) once trained on a setting (such as satellite images), the model is able to discover the changes correlating monotonically with time in unseen image sequences in the same setting, and we demonstrate several situations where this can be applied. Finally, (iv) we demonstrate that the trained model is able to order novel image sequences surpassing the performance of previous approaches for ordering on standard benchmarks.
@@ 3.3 Training and inference
Temporal loss. Given the ground-truth order $y\in\mathbb{Z}^{Q}:y_{i}\in\{0,1,...,F-1\}$, the model can be trained via binary cross-entropy loss. Specifically, we convert the ground-truth order into a binary permutation matrix. With some notation abuse, we still call this matrix $y\in\mathbb{Z}^{F\times Q}$. The forward loss is then simply the elementwise binary cross-entropy between the two matrices: $\mathcal{L}_{f}(y,\hat{y})=\frac{1}{FQ}\sum_{i\in F}\sum_{j\in Q}\text{cross-entropy}(\hat{y}_{ij},y_{ij})$.
In practice, we find that allowing reversibility in the loss aids with training, as many changes are reversible in nature without prior knowledge of the arrow of time [65] (the sequence could equally be ordered from first to last, or last to first). To allow this, we calculate the loss as the minimum of the loss for both forward and backward sequences, i.e. $\mathcal{L}_{r}=\min(\mathcal{L}_{f}(y,\hat{y}),\mathcal{L}_{f}(\text{reverse}(y),\hat{y}))$. This loss is zero when the model predicts the order correctly in either direction.
@@ 3.4 Discussion
From localization to segmentation. While the attribution map is useful, some applications may benefit from going beyond just localization. Here, we propose three solutions. (i) We can directly obtain segmentation at patch-level granularity by thresholding the attribution map, together with minimal post-processing namely averaging across frames and removing small contours. (ii) We replace the linear projection layer on image patches with a pretrained DINOv2 to enhance the feature quality. (iii) Alternatively, we can use the highest activation points (i.e. centre pixels of patches) as prompt for the pretrined Segment Anything model (SAM) [34], to obtain pixel-level segmentation masks.
@@ 4.1 Video datasets
To leverage the inherent temporal information in videos, we sample a sequence from a video, shuffle the frames, then train the model with the original ordering as groundtruth. This enables the attribution map to identify monotonic changes that contribute to the ordering while disregarding other changes. Our study explores sequences across multiple domains, each with distinct cues of interest, which we summarize in Table 1 and illustrate sample sequences in Figure 3.
@@
Timelapse scenes. We gathered outdoor timelapse videos from WebVid-10M dataset [4] with ‘timelapse’ as the search query. The dataset comprises 180 static videos, and our aim is to investigate the cues that the model can extract from the scene to learn its order, as there are no specific absolute cues present.
@@ 5.1 Results on ordering video frames or image sequence
Video ordering results. The results for ordering as well as the main signals that the model can pick up, are shown by EM and EW in Table 1, with qualitative results in Figure 5. The model is able to successfully order sequences across different datasets, particularly in cases where changes are significant. It is expected that the scores are not perfect, as sampled data from videos is not guaranteed to contain ordering cues, as illustrated in the Supplementary.
@@
Failure cases. A limitation of our model is that we do not force a one-to-one matching between queries and images, and this may result in some images being claimed by multiple different queries or by none at all, as seen in Figure 6(b). This problem can easily be resolved by allowing each image to be predicted once. However, not being able to order also provides valuable information as not all real sequences can be ordered – for example, sequences where everything is static for a period, or very stochastic. Therefore, we treat invalid orderings as a means to provide information on whether particular frames can be ordered or not. We report further experimental analyses on the failure cases in the Supplementary.
@@ 5.2 Comparison with change detection methods
Ours 83.3 37.9 34.2 75.0
Ours + DINOv2 80.0 41.3
Ours + SAM 83.3 45.1
Table 2: Localization and segmentation results via the pointing game accuracy for localization, and mIoU for segmentation. The methods for segmentation (patch and SAM) are described in Section 3.4.
We compare against three previous approaches for the ability to detect monotonic temporal changes on Monotonic MUDS dataset, namely, a supervised Siamese Networks [27] trained on MUDS dataset for the task of urban development tracking, by highlighting the differences between buildings; Change Vector Analysis [43], which is a baseline for the task of change detection, highlighting all changes in an image pair without knowledge their nature, and Deep CVA [55], a learned version of CVA that has been trained specifically on images.
The performance comparison is given in Table 2, and illustrated in Figure 7. As can be seen, the urban development tracking method [27] under-segments changes that are correlated with urbanisation, as it is only trained to look at buildings. The change detection methods of [43, 55] over-segments changes that are non-monotonic as it has no concept of time. Both prior methods have shortcomings in detecting urban development: the former method misses changes like road building and deforestation, and the latter includes many erroneous regions such as the seasonal changes in vegetation and water level. Our model is able to highlight correctly the monotonic changes while being invariant to other changes. We further note that our model discovers such changes without any prior information on what to look for. Additionally, replacing the patch projection with DINOv2 and using the localization to prompt SAM both improve the results. We also evaluate on two other datasets: RDS and MoCA, where we have ground-truth for the moving objects in the video; and show that we obtain favourable results. Note that Siamese and DCVA are only trained on satellite images, hence do not generalize to other domains.
@@
Ours 83.3 45.1 30.2
Table 3: Comparison with self-supervised methods on localization and segmentation on Mono-MUDS and on fine-tuning for building change detection on MUDS.
@@ 5.3 Comparison on self-supervised proxy tasks
Here, we compare to other self-supervised methods based on time. We include a discussion in the Supplementary on the subtle differences between the proxy tasks. We include results for training baselines from scratch on MUDS, and testing on Mono-MUDS in localizing and segmenting monotonic changes. The results are shown in Table 3, where we conduct three sets of experiments, as detailed below.
@@
Third, we investigate fine-tuning on a target task that requires both time and localization: change detection of buildings (like [27]). For each method, we pre-train the encoder on MUDS, freeze it, then train a lightweight head (3×3$\times$3 × deconvolutions) on top of the same dataset, and test on unseen sequences. To keep this fair, we keep the number of parameters roughly the same across methods. The results (Table 3, right column) show that our method learns good representations as compared to previous self-supervised methods in localization tasks.
@@ 5.4 Comparison with image ordering methods
We compare to two previous methods on image ordering benchmarks where the task is to arrange the images in increasing order. Differentiable sorting networks such as DiffSort [52] and its successor DSortv2 [53] employ a parameter-free sorting network to rank scalars. Pointer Networks  [63] ranks features by using a recurrent encoder-decoder network with attention. We note that Ptr-Net is not initially designed for such a task, but for arranging a set of coordinates. We simply extend pointer networks by adding an image encoder and task the model to rank the image features from small to large. We then jointly train this encoder and the pointer network. For fair comparison, we use the same transformer encoder as our model, and use the pointer network as the decoder with similar size to our transformer decoder.
@@ 0.F.5 Choice of objective function
0.F.6 Image ordering baselines
In the main paper, we compare our method against baselines for ordering image sequences. We note that the numbers reported in the paper for differentiable sorting methods use a different backbone compared to ours. We experiment with the same backbone, and the performance remains worse than our method throughout both datasets and all metrics.
MNIST: 5 9 16 SVHN: 5 9 16
DSort 78.9||||89.6 72.9||||90.0 48.9||||85.5||||91.1 34.5||||63.6 8.7||||53.0 0.0||||36.1||||40.4
DSv2 85.8||||91.9 76.5||||90.2 60.7||||89.0||||91.6 43.2||||69.9 16.0||||60.3 0.0||||45.6||||54.1
Ours 93.9||||96.7 87.9||||95.2 72.2||||91.2||||92.9 77.3||||88.2 53.9||||81.0 19.4||||67.9||||67.6
Table 14: Backbone ablation. We show that our method remains superior at ordering despite using the same backbone.